# The Endless Tuning An Artificial Intelligence Design To Avoid Human Replacement and Trace Back Responsibilities

### v1

@@ 3 To the thing itself: the protocol
[59] defined Hybrid Intelligence as «the ability to achieve complex goals by combining human and artificial intelligence, thereby reaching superior results to those each of them could have accomplished separately, and continuously improve by learning from each other». [60] noted that using interpretability to increase complementary team performance (CTP) might be a tricky issue: findings may be various. Ex multis, experimenting with various conditions across two risk assessment tasks, [61] found that asking users to first make their own decision led to the highest CTP, surpassing even XAI assistance. [6] experimentally found that XAI should be used adaptively, depending on the model’s confidence. In addition, one issue (see f.e. [62]) is that the user’s solo performance is not always superior to that of the model with which they collaborate (although discovering patterns can enhance human performance [63]). This raises an important reflection: granted that both the human and the model should be as accurate as possible, if it is true that an accurate and “persuasive” model can lead to good outcomes, what about responsibility for errors? Put differently, if we had to choose, would it be preferable to pursue higher accuracy with less control over potential negative consequences, or slightly lower accuracy with greater capacity for remediation? We should also take into account the context in which the task is framed as well as, strictly speaking, the design of the interface itself. Is it useful, for instance, to display the model’s confidence? [7] found that it is; however, users may even “rebel” against the recommendations [64]. After all, we are dealing with the human being, the quintessential “polytropos”.
@@
An application protocol was developed for this study. As for relational approaches and protocols which investigate in a similar direction, we should mention here [65, 66] and above all [67] which has been inspiring with regard to some technical solution. Our protocol consists of five components and is modular, meaning components may be individually excluded depending on specific needs. It takes the form of a dialogue with the artificial intelligent system, structured as a double negative feedback loop that can be navigated both back and forth and in a continuous flow. A multi-step interface has been designed (see Figure 2), through which the user, guided by a series of prompts and given the option to change each time idea, arrives at a potential final decision regarding a given case (without imposing any time limitation), respectively: first impression, explanation as suggestion, similarity as suggestion, confidence measures, and finetuning.
@@ 3.1 What’s your first impression?
A second nuance concerns the very notion of “beginning”. This first step is something of a false start — or better, an aporetic beginning [17, pp. 73-76]: at minimum the third, fourth, or fifth such beginning, following the initial data collection, its labeling, the development of the model, and subsequently the model’s automated training. Such “aporia of the beginning” already signals the reality that our relationship with an artificial intelligence system is, in many ways, endless — both vertically, as it may unfold in an infinite loop, and horizontally, given the length and complexity of the supply chain embedded within the product we are handling.
@@ 3.2 I’ll give you a suggestion: explanations, with a hermeneutic approach
As a premise, in [2] this step was originally placed third, following the similarity comparison. However, the earliest trials of the interface, informally conducted, revealed that users, perhaps due to some “cognitive laziness of their System 2” [4], tended to skip the similarity cues altogether, instead hastily placing significant weight on the allure of saliency maps. As a consequence, this sequence was experimentally inverted, which proved to be more effective.
@@
At this stage, the model has already produced an outcome, but this second step involves presenting the user either with an account of the process that allowed the model to reach its conclusion (e.g., rule-based explanations for decision trees) or with indicators of potential significance w.r.t. input data (i.e., saliency maps), but before knowing the outcome of the model, avoiding “apophantic” interfaces, due to some reasons. Firstly, in the wild, it cannot be taken for granted that the user knows the ground truth. Secondly, an explanation is a double jump: it is not relative to the input data, but to how the machine saw it. On one side, there is no absolute guarantee that the model’s prediction is correct; on the other side, the explanation itself might be unfaithful. Moreover, if the model “reasoned” well (and explanations might be managed, see f.e. [68]) it can provide useful suggestions; whereas, should it have reasoned badly, the user should not yet be so affected by automation bias and nonetheless be able to find out some interesting patterns, some anomalies, as telling them: «hey, look here and there».
@@
A hermeneutics of artificial intelligence is carried out in nuce, using explanations as interpretations. However, are such a term and such a choice justifiable? After all, once the desire to avoid the replacement of the user by the machine has been established, one might argue that there is no need to go so far. After all, the local rules extracted from a decision tree are faithful by design and quantitative metrics such as stability and fidelity exist to assess black boxes’ explanations [69]. At first glance, then, there does not appear to be that necessary distance between the signifier and the signified which typically characterizes such an approach (see f.e. [70] for a survey). Nevertheless, approaching so-called black-boxes, the very notion of a “local explanation” appears somewhat oxymoronic. An explanation, indeed, often bears the character of a strong hypothesis and thus, if not itself a general causal nexus, is at least derived from a general assumption. Instead, for example, the subsymbolic and connectionist functioning of neural networks is akin to what Aristotle referred to as symbebekos, the “concomitant”, which cannot be elevated to logos, even though, in this case, it responds to a precise function. Nonetheless, to explain means to explain for us.
@@
On the one hand, Heidegger [71, §32] already pointed out how our interpretation, including scientific knowledge, always operates within a pre-comprehension, a horizon of meaning, even constituting a vicious circle. Such pre-comprehension cannot directly grasp the inner workings of the algorithmic process and, furthermore, could be strongly influenced by a series of variables that escape the algorithm’s processing. On the other hand, might the XAI explainers be qualified as second-order interpreters? A sign, Charles Peirce wrote [72, 2.228],
@@
Now, as an example, let us take the INTGRAD algorithm [73]. It generates a local explanation by progressively sending a spectrum from a baseline to the chosen input (say, the sign) to a black-box (global function: first piece of the ground), calculating the path integral (chosen process: second piece of the ground) with respect to the features, and thereby extracting their significance. Since the space $x-x^{\prime}$ only represents a small portion of the entire possible input space, it seems that the explanation will inevitably be affected by a bit of vagueness. Still, to further expand such point, the fact remains that multiple valid explanations are possible (see [74] on the “multiplicity of good models”), and we are not only referring to forms of instability caused by interpretation attacks (that is, when two slightly different input instances produce very different explanations). Rather, even when equally good metrics are achieved, if (as will be shown in Section 5) an explainer like RISE [75] observes the behavior of an entire neural network, while Grad-CAM [76] observes only a single layer thereof, why should we expect them to provide the same explanation? In this play of “one, none, and a hundred thousand” black boxes, it seems appropriate to explicitly highlight some distance between signifier and signified. Not to mention the field of textual explanations, to which hermeneutics would be naturally suited, the very concept of a saliency map is intriguing. Its “filagree” (or perhaps “watermark” might be a better term) is bounded both by computational constraints (e.g. the number of masks for RISE) and by the peculiarities of our vision, which is sensitive to color and, above all, low-dimensional. In his Metaphysical Meditations, Descartes would wrote:
«When I imagine a triangle, I do not conceive it only as a figure comprehended by three lines, but I also apprehend these three lines as present by the power and inward vision of my mind, 24 and this is what I call imagining. But if I desire to think of a chiliagon […], I certainly conceive truly that it is a figure composed of a thousand sides […]; but I cannot in any way imagine the thousand sides of a chiliagon, nor do I, so to speak, regard them as present [with the eyes of my mind]»[77, p. 25].
@@ 3.3 I’ll give you a suggestion: similarity
Drawing from Peirce’s concept of thirdness, [78] describes similarity in both physical and epistemological terms as a halfway relationship — as if positioned along a branch of a hyperbola — between identity and difference. So as to facilitate a more or less indirect dialogue between the user and the experts who compiled the training dataset, the possibility is provided of comparing the current case study with the most similar ones in the original dataset, using the original labels. On the one hand, mapping the training data space tends to shift back in time the center of responsibility, particularly if not placing conditions on the labels. On the other hand, putting a bit more power back in the algorithm’s hands, clustering algorithms might be applied, for example on a latent space. Yet, as experience taught us, it is a matter of subtle and tricky balance.
An objection may arise, already raised by [78]: can similarity be computed? After all, it has a formal nature, whereas we cannot help deeming it a geometric proximity measure. However, not only does the activity of a subject (the data annotator) not seem absent at all, but nothing prevents f.e. an autoencoder computing embeddings to reserve surprises in its latent space. In fact, to us, this last one seemed also the best way to account for difference, because it can introduce a margin of variance. Actually, differences usually emerge through criticisms (the more distant cases from the same class) and counterfactuals (the more similar cases from a different class). Such features are useful in the context of explainability, and at first glance, it might seem that the user’s critical thinking could benefit from them. However, here we do not have to explain the user’s choice as the one of a black box, but rather guide them through their decisions. The more, using criticisms or counterfactuals, which label should be taken as the basis? Since we cannot use the label provided by the classifier model (because it is still “secret”), we would need to use the last label provided by the user. Such a choice would entail a risk. If we told the user, e.g., something like: «look at this other case now: it lies quite far from the one you are observing, yet it has the same label», if the their opinion were correct, they would be misled.
@@ 3.4 Do you really want to know my opinion?
A model never explicitly says “yes” or “no”; rather, it provides an indication. A typical case is regression. Nonetheless, particularly in classification settings, this indication might be structured, often through a softmax function applied to the last layer of a neural network, as a probability measure for the various options, thus functioning as a measure of confidence. We could broadly define confidence as the model’s certainty “in itself” concerning its own prediction, in contrast to its actual output (if it were solely about the latter, we might perhaps more appropriately refer to it as accuracy). This module, the penultimate of the protocol and the last of the interface, consists of presenting the user with the actual outcome of the model, which has been hidden up to this point, preferably together with its confidence. The purpose of providing the result alongside the confidence, as noted by the aforementioned [7], is to assist the user in calibrating their trust in the device (after all, this is a strategy both against automation bias and algorithmic aversion), while, in full alignment with the relational ethical approach proposed here, the human is required to contribute their own knowledge. However, caution is needed with overconfident systems (see f.e. [79] for relative issues). On the one hand, therefore, it would be useful to warn the user; on the other hand, paraphrasing [15], isn’t the soft voice of Prudence more ear-friendly than the one of Bluster?
@@ 3.5 Finetuning
In addition, it is necessary to underline at this point a socio-technical limitation in this protocol (which is in fact a limitation of artificial intelligence itself): the need for clean and robust data. Due to such a limitation, for example, it might happen that a single user will not be able to train and fine-tune a model through such a protocol by themselves alone. Too many work sessions could be needed, or a personal computer could not have the minimum computational power available, implying the need for a remote machine which in turn would entail more privacy issues. Almost constituting an “environmental” expansion of the ethical approach pursued here, we would like to place our hope (precisely due to the lightness of the continuous learning approach) in the creation of “socio-technical clusters” — for example, asking a team of doctors from the same hospital (or at most, doctors belonging to the same geographical area of healthcare) to join their efforts in a collaborative perspective as encouraged by the European Data Act [80], seeking to benefit from locally hosted servers and getting engaged in the robustification of the data.
@@ 4.1 Prototypes
The first version concerns the task of granting or denying a loan to a fictitious applicant, according to some indices. As a consequence, it is a binary classification task. We used the Loan Approval Classification Dataset [81] made available on Kaggle, inspired by [82] and [83], and licensed under the Apache 2.0 License. The dataset consisted of a .csv file containing 45,000 applicants and 13 attributes: age, gender, education, income, employment experience, home ownership, loan amount, loan intent, loan interest rate, loan percent income, credit history length, credit score, and previous loan defaults (plus the target: grant or deny). To obtain a perfectly balanced dataset, exceeding items were randomly removed, obtaining a training dataset of 18,000 rows. At the same time, a balanced “case study set” of 200 items was extracted, together with a “temporary set” of 1795 items used as a memory, so to say a “piggy bank”. The 200 case studies were available to be chosen by the expert during the test. A decision tree was trained on the training set, reaching an accuracy of $\approx 83\%$. Since our aim was simply to provide a reasonably accurate model, in all the three cases the case study set was employed as a test set, on which accuracy reached even a slightly higher percentage.
@@
In the second case, we made use of the WikiArt dataset [84], made available on Kaggle with CC0 License, and originally sourced from www.wikiart.org. The dataset initially consisted of approximately 80,000 images, spanning 27 different artistic styles. For class balance purposes, the number of styles considered in the task was reduced to seven (those with the highest number of images): art nouveau, baroque, impressionism, expressionism, symbolism, post-impressionism, romanticism. Out of the remaining $\approx 28,000$ images, which constituted our training set, 70 were selected as case studies, while 420 images composed the temporary set. A 34-layer ResNet [85], pretrained on ImageNet1K, was therefore fine-tuned on $\approx 27{,}500$ images, achieving an accuracy of $\approx 65\%$. All the images were resized to 1080x1080 px, and black padding was added so as to compensate for asymmetries.
@@
The third task was a binary classification problem aimed at diagnosing the presence or absence of pneumonia in chest X-ray images. In this case, we employed the Chest X-Ray Images (Pneumonia) dataset [86], made available on Kaggle under the CC BY 4.0 license, and inspired by [87, 88]. Originally composed of $\approx 5800$ files, due to class imbalance, the dataset was randomly reduced to 2500 items as a training set, according to the labels pneumonia: yes and pneumonia: no. Moreover, the case study set consisted of 100 items, while the temporary set consisted of 198 items. Also here, resizing to 1080x1080 px together with black padding were carried out. Another 34-layer Resnet, identical to the first one, was finetuned on such a training set, reaching a test accuracy of $\approx 95\%$. A couple of further layers were added to both the networks: a linear layer with output_size=14, and a final linear layer with output_size=num_classes. In both the second and third experiment, black padding was felt necessary because the recognition also depended on the shape of the objects depicted. It was preferred to give the right of way to the expert, making the models train on padded images, rather than to dangerously distort the data.
@@
In the first case, in form of decision rules extracted from the tree in a natural language fashion and variously colored — not casually, our tree was bounded to max_depth=4, so as to avoid confusion (see Figure 3(a) for an example). Such rules are obviously truthful. Text strings were color-coded as a form of cognitive reinforcement, with careful attention to using equally vivid colors that would not be interpreted in a conventional way — such as green for “grant” or bright red for “deny”. In the other two cases, drawing on the hermeneutic differences discussed in Section 3.2, the user is presented with two separate pages, each displaying a distinct saliency map. The first is generated using RISE [75] (slightly optimized due to computational constraints, producing 1,500 masks), while the second is obtained through Grad-CAM [89, 76]), applied to the GAP layer. These two visualizations reflect different perspectives, on the side of the explainers, on the neural network’s decision-making process. Both maps were carried out using the “jet” colormap, going from deep blue to intense red.

### v3

@@ 3 To the thing itself: the protocol
[dellermann2019hybrid] defined Hybrid Intelligence as «the ability to achieve complex goals by combining human and artificial intelligence, thereby reaching superior results to those each of them could have accomplished separately, and continuously improve by learning from each other». [hemmer2021human] noted that using interpretability to increase complementary team performance (CTP) might be a tricky issue: findings may be various. Ex multis, experimenting with various conditions across two risk assessment tasks, [green2019principles] found that asking users to first make their own decision led to the highest CTP, surpassing even XAI assistance. [bansal2021does] experimentally found that XAI should be used adaptively, depending on the model’s confidence. In addition, one issue (see f.e. [alufaisan2021does]) is that the user’s solo performance is not always superior to that of the model with which they collaborate (although discovering patterns can enhance human performance [lai2020chicago]). This raises an important reflection: granted that both the human and the model should be as accurate as possible, if it is true that an accurate and “persuasive” model can lead to good outcomes, what about responsibility for errors? Put differently, if we had to choose, would it be preferable to pursue higher accuracy with less control over potential negative consequences, or slightly lower accuracy with greater capacity for remediation? We should also take into account the context in which the task is framed as well as, strictly speaking, the design of the interface itself. Is it useful, for instance, to display the model’s confidence? [zhang2020effect] found that it is; however, users may even “rebel” against the recommendations [dietvorst2015algorithm]. After all, we are dealing with the human being, the quintessential “polytropos”.
An application protocol was developed for this study. As for relational approaches and protocols which investigate in a similar direction, we should mention here [cabitza2022open, cabitza2023rams] and above all [campisi2023doctoral] which has been inspiring with regard to some technical solution. Moreover, in line with Miller’s proposal of an evaluative AI framework [miller2023explainable], we acknowledge the value of a machine-in-the-loop approach that privileges abductive reasoning and user reflection over mere recommendation. We concur that explanations should stimulate hypothesis formation rather than provide definitive answers. Nevertheless, certain aspects could call for deeper examination. The fact is that psychology and ethics do not necessarily align, and under conditions of stress or cognitive fragility, a form of near-recommendation may prove legitimate (what, to tell the truth, the evaluative AI framework would be ready to, being sufficiently elastic). Once an AI system assumes a prompting role, questions of opacity, responsibility, and mutual learning from error might become increasingly complex. For exmple, the possible assumption that evaluative AI could remain value-neutral would appear questionable: if models are trained on biased data, we ask whether their explanations inevitably reflect such biases. More broadly, one may conclude that models themselves possess a degree of agency — a character of interpreters whose epistemic and ethical weight should perhaps be explicitly acknowledged.
@@
Our protocol consists of five components and is modular, meaning components may be individually excluded depending on specific needs. It takes the form of a dialogue with the artificial intelligent system, structured as a double negative feedback loop that can be navigated both back and forth and in a continuous flow. A multi-step interface has been designed (see Figure 2), through which the user, guided by a series of prompts and given the option to change each time idea, arrives at a potential final decision regarding a given case (without imposing any time limitation), respectively: first impression, explanation as suggestion, similarity as suggestion, confidence measures, and finetuning.
@@ 3.1 What’s your first impression?
A second nuance concerns the very notion of “beginning”. This first step is something of a false start — or better, an aporetic beginning [fabris2016relazione, pp. 73-76]: at minimum the third, fourth, or fifth such beginning, following the initial data collection, its labeling, the development of the model, and subsequently the model’s automated training. Such “aporia of the beginning” already signals the reality that our relationship with an artificial intelligence system is, in many ways, endless — both vertically, as it may unfold in an infinite loop, and horizontally, given the length and complexity of the supply chain embedded within the product we are handling.
@@ 3.2 I’ll give you a suggestion: explanations, with a hermeneutic approach
As a premise, in [fabris2024towards] this step was originally placed third, following the similarity comparison. However, the earliest trials of the interface, informally conducted, revealed that users, perhaps due to some “cognitive laziness of their System 2” [kahneman2011thinking], tended to skip the similarity cues altogether, instead hastily placing significant weight on the allure of saliency maps. As a consequence, this sequence was experimentally inverted, which proved to be more effective.
@@
At this stage, the model has already produced an outcome, but this second step involves presenting the user either with an account of the process that allowed the model to reach its conclusion (e.g., rule-based explanations for decision trees) or with indicators of potential significance w.r.t. input data (i.e., saliency maps), but before knowing the outcome of the model, avoiding “apophantic” interfaces, due to some reasons. Firstly, in the wild, it cannot be taken for granted that the user knows the ground truth. Secondly, an explanation is a double jump: it is not relative to the input data, but to how the machine saw it. On one side, there is no absolute guarantee that the model’s prediction is correct; on the other side, the explanation itself might be unfaithful. Moreover, if the model “reasoned” well (and explanations might be managed, see f.e. [teso2019explanatory]) it can provide useful suggestions; whereas, should it have reasoned badly, the user should not yet be so affected by automation bias and nonetheless be able to find out some interesting patterns, some anomalies, as telling them: «hey, look here and there».
@@
A hermeneutics of artificial intelligence is carried out in nuce, using explanations as interpretations. However, are such a term and such a choice justifiable? After all, once the desire to avoid the replacement of the user by the machine has been established, one might argue that there is no need to go so far. After all, the local rules extracted from a decision tree are faithful by design and quantitative metrics such as stability and fidelity exist to assess black boxes’ explanations [bodria2023benchmarking]. At first glance, then, there does not appear to be that necessary distance between the signifier and the signified which typically characterizes such an approach (see f.e. [bianco2007introduzione] for a survey). Nevertheless, approaching so-called black-boxes, the very notion of a “local explanation” appears somewhat oxymoronic. An explanation, indeed, often bears the character of a strong hypothesis and thus, if not itself a general causal nexus, is at least derived from a general assumption. Instead, for example, the subsymbolic and connectionist functioning of neural networks is akin to what Aristotle referred to as symbebekos, the “concomitant”, which cannot be elevated to logos, even though, in this case, it responds to a precise function. Nonetheless, to explain means to explain for us.
@@
On the one hand, Heidegger [heidegger1962heidegger, §32] already pointed out how our interpretation, including scientific knowledge, always operates within a pre-comprehension, a horizon of meaning, even constituting a vicious circle. Such pre-comprehension cannot directly grasp the inner workings of the algorithmic process and, furthermore, could be strongly influenced by a series of variables that escape the algorithm’s processing. On the other hand, might the XAI explainers be qualified as second-order interpreters? A sign, Charles Peirce wrote [peirce1931collected, 2.228],
@@
Now, as an example, let us take the INTGRAD algorithm [sundararajan2017axiomatic]. It generates a local explanation by progressively sending a spectrum from a baseline to the chosen input (say, the sign) to a black-box (global function: first piece of the ground), calculating the path integral (chosen process: second piece of the ground) with respect to the features, and thereby extracting their significance. Since the space $x-x^{\prime}$ only represents a small portion of the entire possible input space, it seems that the explanation will inevitably be affected by a bit of vagueness. Still, to further expand such point, the fact remains that multiple valid explanations are possible (see [breiman2001statistical] on the “multiplicity of good models”), and we are not only referring to forms of instability caused by interpretation attacks (that is, when two slightly different input instances produce very different explanations). Rather, even when equally good metrics are achieved, if (as will be shown in Section 5) an explainer like RISE [petsiuk2018rise] observes the behavior of an entire neural network, while Grad-CAM [selvaraju2017grad] observes only a single layer thereof, why should we expect them to provide the same explanation? In this play of “one, none, and a hundred thousand” black boxes, it seems appropriate to explicitly highlight some distance between signifier and signified. Not to mention the field of textual explanations, to which hermeneutics would be naturally suited, the very concept of a saliency map is intriguing. Its “filagree” (or perhaps “watermark” might be a better term) is bounded both by computational constraints (e.g. the number of masks for RISE) and by the peculiarities of our vision, which is sensitive to color and, above all, low-dimensional. In his Metaphysical Meditations, Descartes would wrote:
«When I imagine a triangle, I do not conceive it only as a figure comprehended by three lines, but I also apprehend these three lines as present by the power and inward vision of my mind, 24 and this is what I call imagining. But if I desire to think of a chiliagon […], I certainly conceive truly that it is a figure composed of a thousand sides […]; but I cannot in any way imagine the thousand sides of a chiliagon, nor do I, so to speak, regard them as present [with the eyes of my mind]»[descartes_meditations, p. 25].
@@ 3.3 I’ll give you a suggestion: similarity
Drawing from Peirce’s concept of thirdness, [marcos2011sem] describes similarity in both physical and epistemological terms as a halfway relationship — as if positioned along a branch of a hyperbola — between identity and difference. So as to facilitate a more or less indirect dialogue between the user and the experts who compiled the training dataset, the possibility is provided of comparing the current case study with the most similar ones in the original dataset, using the original labels. On the one hand, mapping the training data space tends to shift back in time the center of responsibility, particularly if not placing conditions on the labels. On the other hand, putting a bit more power back in the algorithm’s hands, clustering algorithms might be applied, for example on a latent space. Yet, as experience taught us, it is a matter of subtle and tricky balance.
An objection may arise, already raised by [marcos2011sem]: can similarity be computed? After all, it has a formal nature, whereas we cannot help deeming it a geometric proximity measure. However, not only does the activity of a subject (the data annotator) not seem absent at all, but nothing prevents f.e. an autoencoder computing embeddings to reserve surprises in its latent space. In fact, to us, this last one seemed also the best way to account for difference, because it can introduce a margin of variance. Actually, differences usually emerge through criticisms (the more distant cases from the same class) and counterfactuals (the more similar cases from a different class). Such features are useful in the context of explainability, and at first glance, it might seem that the user’s critical thinking could benefit from them. However, here we do not have to explain the user’s choice as the one of a black box, but rather guide them through their decisions. The more, using criticisms or counterfactuals, which label should be taken as the basis? Since we cannot use the label provided by the classifier model (because it is still “secret”), we would need to use the last label provided by the user. Such a choice would entail a risk. If we told the user, e.g., something like: «look at this other case now: it lies quite far from the one you are observing, yet it has the same label», if the their opinion were correct, they would be misled.
@@ 3.4 Do you really want to know my opinion?
A model never explicitly says “yes” or “no”; rather, it provides an indication. A typical case is regression. Nonetheless, particularly in classification settings, this indication might be structured, often through a softmax function applied to the last layer of a neural network, as a probability measure for the various options, thus functioning as a measure of confidence. We could broadly define confidence as the model’s certainty “in itself” concerning its own prediction, in contrast to its actual output (if it were solely about the latter, we might perhaps more appropriately refer to it as accuracy). This module, the penultimate of the protocol and the last of the interface, consists of presenting the user with the actual outcome of the model, which has been hidden up to this point, preferably together with its confidence. The purpose of providing the result alongside the confidence, as noted by the aforementioned [zhang2020effect], is to assist the user in calibrating their trust in the device (after all, this is a strategy both against automation bias and algorithmic aversion), while, in full alignment with the relational ethical approach proposed here, the human is required to contribute their own knowledge. However, caution is needed with overconfident systems (see f.e. [zhang2023survey] for relative issues). On the one hand, therefore, it would be useful to warn the user; on the other hand, paraphrasing [miller2023explainable], isn’t the soft voice of Prudence more ear-friendly than the one of Bluster?
@@ 3.5 Finetuning
In addition, it is necessary to underline at this point a socio-technical limitation in this protocol (which is in fact a limitation of artificial intelligence itself): the need for clean and robust data. Due to such a limitation, for example, it might happen that a single user will not be able to train and fine-tune a model through such a protocol by themselves alone. Too many work sessions could be needed, or a personal computer could not have the minimum computational power available, implying the need for a remote machine which in turn would entail more privacy issues. Almost constituting an “environmental” expansion of the ethical approach pursued here, we would like to place our hope (precisely due to the lightness of the continuous learning approach) in the creation of “socio-technical clusters” — for example, asking a team of doctors from the same hospital (or at most, doctors belonging to the same geographical area of healthcare) to join their efforts in a collaborative perspective as encouraged by the European Data Act [EU2023DataAct], seeking to benefit from locally hosted servers and getting engaged in the robustification of the data.
@@ 4.1 Prototypes
The first version concerns the task of granting or denying a loan to a fictitious applicant, according to some indices. As a consequence, it is a binary classification task. We used the Loan Approval Classification Dataset [tai_loan_dataset] made available on Kaggle, inspired by [credit_risk_dataset] and [financial_risk_dataset], and licensed under the Apache 2.0 License. The dataset consisted of a .csv file containing 45,000 applicants and 13 attributes: age, gender, education, income, employment experience, home ownership, loan amount, loan intent, loan interest rate, loan percent income, credit history length, credit score, and previous loan defaults (plus the target: grant or deny). To obtain a perfectly balanced dataset, exceeding items were randomly removed, obtaining a training dataset of 18,000 rows. At the same time, a balanced “case study set” of 200 items was extracted, together with a “temporary set” of 1795 items used as a memory, so to say a “piggy bank”. The 200 case studies were available to be chosen by the expert during the test. A decision tree was trained on the training set, reaching an accuracy of $\approx 83\%$. Since our aim was simply to provide a reasonably accurate model, in all the three cases the case study set was employed as a test set, on which accuracy reached even a slightly higher percentage.
@@
In the second case, we made use of the WikiArt dataset [wikiart_dataset], made available on Kaggle with CC0 License, and originally sourced from www.wikiart.org. The dataset initially consisted of approximately 80,000 images, spanning 27 different artistic styles. For class balance purposes, the number of styles considered in the task was reduced to seven (those with the highest number of images): art nouveau, baroque, impressionism, expressionism, symbolism, post-impressionism, romanticism. Out of the remaining $\approx 28,000$ images, which constituted our training set, 70 were selected as case studies, while 420 images composed the temporary set. A 34-layer ResNet [he2016deep], pretrained on ImageNet1K, was therefore fine-tuned on $\approx 27{,}500$ images, achieving an accuracy of $\approx 65\%$. All the images were resized to 1080x1080 px, and black padding was added so as to compensate for asymmetries.
@@
The third task was a binary classification problem aimed at diagnosing the presence or absence of pneumonia in chest X-ray images. In this case, we employed the Chest X-Ray Images (Pneumonia) dataset [chestxraydataset], made available on Kaggle under the CC BY 4.0 license, and inspired by [kermany2018identifying, kermany2018labeled]. Originally composed of $\approx 5800$ files, due to class imbalance, the dataset was randomly reduced to 2500 items as a training set, according to the labels pneumonia: yes and pneumonia: no. Moreover, the case study set consisted of 100 items, while the temporary set consisted of 198 items. Also here, resizing to 1080x1080 px together with black padding were carried out. Another 34-layer Resnet, identical to the first one, was finetuned on such a training set, reaching a test accuracy of $\approx 95\%$. A couple of further layers were added to both the networks: a linear layer with output_size=14, and a final linear layer with output_size=num_classes. In both the second and third experiment, black padding was felt necessary because the recognition also depended on the shape of the objects depicted. It was preferred to give the right of way to the expert, making the models train on padded images, rather than to dangerously distort the data.
@@
In the first case, in form of decision rules extracted from the tree in a natural language fashion and variously colored — not casually, our tree was bounded to max_depth=4, so as to avoid confusion (see Figure 3(a) for an example). Such rules are obviously truthful. Text strings were color-coded as a form of cognitive reinforcement, with careful attention to using equally vivid colors that would not be interpreted in a conventional way — such as green for “grant” or bright red for “deny”. In the other two cases, drawing on the hermeneutic differences discussed in Section 3.2, the user is presented with two separate pages, each displaying a distinct saliency map. The first is generated using RISE [petsiuk2018rise] (slightly optimized due to computational constraints, producing 1,500 masks), while the second is obtained through Grad-CAM [jacobgilpytorchcam, selvaraju2017grad]), applied to the GAP layer. These two visualizations reflect different perspectives, on the side of the explainers, on the neural network’s decision-making process. Both maps were carried out using the “jet” colormap, going from deep blue to intense red.